Consider a short, 10-meter link, over which a sender can transmit at a rate of 170 bits/sec in both directions. Suppose that packets containing data are 100,000 bits long, and packets containing only control (e.g., ACK or handshaking) are 300 bits long. Assume that N parallel connections each get 1/N of the link bandwidth. Now consider the HTTP protocol, and suppose that each downloaded object is 100 Kbits long and that the initial downloaded object contains 10 referenced objects from the same sender. Would parallel downloads via parallel instances of non-persistent HTTP make sense in this case? Now consider persistent HTTP. Do you expect significant gains over the non-persistent case? Justify and explain your answer
Answer #1

Length of the link = 10 meters

Bandwidth = 170 bits/sec

Size of a data packet = 100,000 bits

Size of a control packet = 300 bits

Size of the downloaded object = 100Kbits

No. of referenced objects = 10

Consider Tp to be the propagation delay between the server and the client. Let dpbe the propagation delay and dt be the transmission delay.

The total time delay for sending and receiving packets is given by:

d = dp (propagation delay) + dt (transmission delay)

Parallel downloads via parallel instances of non-persistent HTTP:

Bandwidth = 170 bits/sec

No. of referenced objects = 10

Bandwidth for each parallel download = 170/10

= 17 bits/sec

During parallel downloads, 10 independent connections are established, and the objects are downloaded simultaneously on these networks. First, a client sends a request for the object. Next, the server processes the request and once the connection is set, the server sends the object in response.

Therefore, the total time required for parallel downloads, is calculated as:

(300/170 + Tp + 300/170 +Tp +300/170+ Tp + 100,000/170 + Tp) + (300/17 + Tp + 300/17 + Tp + 300/170 + Tp + 100,000/17 + Tp)

= ((300+300+300+100,00)/170 + 4Tp) + ((300+300+300+100,00)/17+ 4Tp)

= ((100,900)/170 + 4Tp) + ((100,700)/17 + 4Tp)

= (593 + 4Tp) + (5935 + 4Tp)

= 6528 + 8 Tp seconds

Thus, parallel instances of non-persistent HTTP makes sense in this case.

Let the propogation speed of the medium be 300*106 m/sec.

Then, Tp = 10/(300*106)

= 0.03 micro seconds

Thus, Tp is negligible. This would lead to a similar value of time delay for both persistent HTTP and non-persistent HTTP. Hence, persistent HTTP is not faster than non-persistent HTTP with parallel downloads.